OpenStudy (sedatefrog712):

Medal Part A: Write the expression x2 + 2x - 24 as a product of two linear expressions. Show your work and justify each step. Part B: Rewrite x2 - 8x + 16 as a square of a linear expression i only need help with part A though

OpenStudy (sedatefrog712):

@triciaal

OpenStudy (sedatefrog712):

part a is all i need help with really

OpenStudy (nathan917):

A) During factoring, we want 2 linear expressions to equal the given quadratic expression. Thus, we find that (x+6)(x-4) would equal x^2+2x-24 B) (x-4)^2... Mix a up as much as ya can :).

OpenStudy (sedatefrog712):

ok thanx

OpenStudy (nathan917):

No prob.

OpenStudy (triciaal):

@nathan917 thanks for your input @SedateFrog712 let's do this systematically

OpenStudy (sedatefrog712):

ok

OpenStudy (triciaal):

x2 + 2x - 24 =0 general form a = 1 b = 2 c = -24 a*c = -24 what are the factors of -24 that add up to +2?

OpenStudy (triciaal):

use -12 and +2 x^2 -12x + 2x -24 =0 rewrite to group if necessary and factor each group

OpenStudy (triciaal):

sorry 6 and -4

OpenStudy (sedatefrog712):

its ok

OpenStudy (triciaal):

x^2 + 6x -4x -24 = 0 x(x + 6) -4(x + 6) =0 (x-4)(x +6) =0 x=4; x = -6

OpenStudy (triciaal):

forget x = question only asked for the factors

OpenStudy (sedatefrog712):

o ok it make allot more scene now thank you for explaining it to me

OpenStudy (triciaal):

anyway the idea is that you do have steps to follow rather that mixing and matching to redo each time.

OpenStudy (triciaal):

and you see the signs are important